# Complexity of Einstein-Maxwell-non-minimal coupling $R^{2}F^{2}$: the role of the penalty factor

### v1

@@ 4.2 The non-minimal coupling $q_{2}$
In this section we argue that the non-minimal coupling $q_{2}$ is addressed to the scrambling time of the boundary theory. The non-minimal term acts as a deformation of Einstein-Maxwell theory. The dual of this theory in two dimensions (by dimension reduction and some compactifications) is Sachdev-Ye-Kitaev (SYK) model [^1] We should note that the duality of SYK model and black holes (Jackiew-Teitelboim gravity) is well-established for the $q_{2}=0$, without the non-minimal coupling. With $q_{2}\neq 0$ the dual theory would be a deformed SYK model; nonetheless, at low-curvature regime it reduces to SYK model. In addition, our results about the curvature penalty factor is intact whether $q_{2}=0$ or not. which has a very little scrambling time. By deformation, we are going away from the fast scrambler but this not a simple relationship.
The scrambling time of the bulk theory is independent of the generalization parameter. However, it seems that the scrambling time should depend on the penalty factor in the quantum circuit, because of the fact that the penalty factor in the circuit makes an effect on the scrambling of information. Nonetheless, in the bulk the role of the penalty factor is ignored. In the following we define an effective scrambling time in the bulk theory call it the new scrambling time and show that how it controls CGR.
As (60) shows the action of the particle in (40) affects in the complexity of the boundary theory. However, by the generalized term the metric is scaled [^2] It is worth mentioning that this is a scaling and not the transformation of the coordinates.:
@@
where $g_{xx}$ is not scaled. In this way, it seems that the scaled metric or the generalized term weights different directions unequally. But we should have it in mind that the rescaling does not change the bulk theory and only the definition of the complexity. In other words, the butterfly velocity and the scrambling time would be intact. In this section we introduce a new butterfly velocity and the scrambling time which is due to the action (40) or (61), where the dynamic of the particle determines the complexity. Put it differently, when the metric is scaled by the generalized term in the complexity functional, the new scrambling time would be changed which signaling the unequal weight of the particle motion.
The new butterfly velocity in $d+1$ dimensions in (61) is determined by [42]:
@@
where the derivatives are with respect to $r$. Then, the new butterfly velocity $v_{(scal.)B}$ is scaled as:
@@
Besides the scrambling time (B-15) for the action (40), we introduce the new scrambling time which is scaled as:
@@
It is believed that there is an inverse relation between the scrambling time and the complexity growth rate, it means that by an increase in the scrambling time, the processing rate would be decreased. However, there is not a universal relation, it is model-dependent. The leading term in (64) would be:
@@
$\approx\frac{1}{t_{scal.}}=\frac{|a(R)|}{t_{*}}$ (74)
@@
The leading order in CGR based on (73), (85), (86) and setting $\gamma=-0.002$ would be:
@@
Fig.s (5b) and (5d) illustrate that CGR is reduced by the increase in the non-minimal coupling.
In this section we showed that the new scrambling time is related to the non-minimal coupling $q_{2}$ and consequently, by the change of the coupling, CGR is changed as is expected.
@@ 4.3 The conserved charge $Q$
In this section we argue that the value of the conserved charge $Q$ restricts the simple unitary operators on the boundary, as result, it restricts the complexity and CGR.
@@
This lemma brings a restriction on the choice of the simple unitary operators. Due to the fact that the accuracy $\epsilon$ should lie in the interval $\Big(0,max[2\sin(|E|/2)]\Big]$, it demonstrates that the eigenvalue of the operator $U$ could affect the choice of the quantum gates.
In this manner, the distance for a particular accuracy is restricted by the value of the conserved charge $Q$, which means that the change in the value of the charge restricts the simple unitary operators and as consequence CGR is restricted.
However, it could be shown approximately that the new scrambling time for the solution (12) has an approximate relation with the conserved charge. Then, by the increase of the value of the conserved charge, the new scrambling time would be increased and as a result the rate of the complexity as the rate of the processing would be decreased.
@@
For $b=C^{2}$: The complexity growth rate based on (73), (75) and (76) is given by:
@@
Fig.s (3b) and (3d) displays that by increase in the conserved charge, CGR is reduced.
@@
Fig.s (5b) and (5d) depict the decrease of CGR by the increase in the conserved charge.
All in all, CGR is reduced by the increase of the conserved charge as is expected of the limitation of the quantum gates in SKA.
4.4 Superconducting qubits circuit
In this section, we provide two examples in superconducting qubits circuits where by an increase in the penalty factor, CGR would be increased or decreased, which is dependent to the penalty factor. The first example shows a direct relation between the penalty factor and CGR: a quantum circuit that CGR is increased by the increasing the penalizing the circuit. Moreover, the second one provides an inverse relation.
In the superconducting qubit circuits to prepare a qubit one needs controlling quickly the fast gates (or low runtime) to stay away from decoherence. However, fast controlling pose a problem because of the fact that it requires a strong, sharp microwave pulse, but strong pulse has a wide frequency spectrum. A strong pulse allows for very fast $CNOT$ gates [45]. Nonetheless, this makes parallelizing operations to reduce circuit depth. Moreover, a wide frequency pulse excite the qubit to higher energy states which leads to the leakage error or crosstalk between energy levels. To reduce the leakage it requires the qubit be more anharmonic that means the energy gap of the states be more significantly different [45]. The anharmonicity makes the states more isolated from the states with higher energies. A problem back of the other, the anharmonicity makes the qubit more sensitive to charge noise which results in reducing the coherence time. As a way out, one needs increase the circuit depth by applying crosstalk penalty.
@@ 4.4 Superconducting qubits circuit
Suppose that there are three qubits of QB1 connected to QB2 and QB2 connected to QB3 where crosstalk is high between QB1&\&QB2 and QB2&\&QB3. If one needs to apply a CNOT(1,3) operator, it is not possible due to the fact that they are not adjacent. Then, it should be applied a SWAP(2,3) operation and then applies CNOT(1,3). However, SWAP(2,3) is not a primitive gate. SWAP is complied into three CNOT operators as the figure (7) shows. Similar occasion is appeared in IBM superconducting quantum circuit [46].
@@
Another example that demonstrates an inverse relation between the penalizing and CGR is in the quantum approximation optimization algorithm (QAOA) where to optimize the algorithm inserts some constraints (as a penalty factor). These constraints make the algorithm more efficient and reduce the circuit depth as an optimization. These algorithm are used in superconducting circuits that are prone to noises and require shallower circuits to prevent noise decohrence. If the evolution of the problem is described by problem Hamiltonian or the system Hamiltonian, then the constraints as an optimization is added to system:
@@
where $\gamma_{a}$ parameter tune the intense of the optimization or penalizing (similar to $\gamma$ in the generalized complexity). Then, the optimization would be the qubit that has the most overlap with the lower eigenstate of the system after $p$ layers of operations or depth $p$. [47] shows such a relation where the intensifying $\gamma_{a}$ makes $p$ layers reduced. It means that by intensifying the obeying rules, the system in shorter steps gets to the desired solution.
By the way, in [48], the authors provide an algorithm known as mixer-phaser ansatz in superconducting processors in which they did not used a penalty-based algorithm. In a layman’s term, if the problem system goes away from your desired solution, the penalty factor penalizes the system to obey the constraints. However, in [48] the constraints have been inserted before the system circuit. Then, they get rid of the penalizing.
The analogue of the superconducting circuits is SYK model that is a solvable model of strange metal states where it has a non-local Hamiltonian [49]. Now, instead of qubits, we are dealing with some strongly interacting fermions, then the Hilbert space of fermions in SYK model could be mapped to the Hilbert space of qubits by Bravyi-Kitaev transformation [50] and Jordan-Wigner transformation [51]. Put it differently, these examples could be transformed into strange metal circuits with fermions.
@@ 5 Conclusion
In this study, we presented a black brane solution in four-dimensional AdS spacetime Einstein-Maxwell theory which is characterized by a non-minimal term $R^{2}F_{\mu\nu}F^{\mu\nu}$. As this model lacks a non-perturbative solution, we addressed it using a perturbative approach based on the non-minimal coupling $q_{2}$ up to first order. This non-minimal coupling can partially describe strange metals; specifically, our model exhibits a linear relationship between the resistivity and the temperature, though it does not include a phase transition. As a consequence, this model could be used to implement quantum computations. Although, superconducting processors are widely used in quantum computations, there are some transformations that make a connection between the qubits of the superconducting circuits and the fermions of the strange metal circuits such as Bravyi-Kitaev and Jordan-Wigner transformations.
We considered the complexity of this model via the complexity equals anything proposal. In this proposal there is an arbitrariness of the generalization parameter $b$. The computations have been done for three choice parameters of $C^{2}$, the Weyl squared tensor, $R^{2}F^{2}$, the squared of the Ricci scalar and the Maxwell field strength, and $F^{2}$. CGR in this model is characterized by three parameters of the conserved charge $Q$, the non-minimal coupling $q_{2}$ and the generalized parameter $b$. We showed that:
  • The generalization parameter acts as a penalty factor in the bulk.
  • The non-minimal coupling is addressed to the effective scrambling time that contains the generalization parameter.
  • The conserved charge is related to Solovay-Kitaev algorithm that states the increase in the value of the conserved charge makes the available simple unitary operators restricted and as consequence the restriction of the processing rate or CGR.
It has been seen that the three generalization parameters have different behavior in change of the non-minimal coupling. CGR in $C^{2}$ and $F^{2}$ is decreased by the $q_{2}$ increase and $R^{2}F^{2}$ by $q_{2}$ diminishment. In other words, a direct and an inverse relationship. Then, we provided two quantum circuit examples in the superconducting quantum circuits that show the behavior of CGR in our solution. These examples are widely applied in superconducting-based computations and demonstrate how intensifying the penalty factor in the circuits could suppressed or encouraged CGR. Moreover, we have noticed that CGR with the generalization parameter $F^{2}$ does not contain branches and its behavior is similar to CV and CA.
In addition, in quantum algorithm has been introduced some algorithms that by removing the penalty factor received more efficient circuits with shorter depth circuits. For future research we think that this could inspire to introduce a holographic complexity proposal that is free of penalizing and restricts the target states instead of penalizing the simple unitary operators. Furthermore, it seems that the generalization parameter of the kind of spacetime quantities like $R$, $R^{2}$ and $C^{2}$ lead to different branches in CGR. For future research this point could be scrutinized that why there is no branch in CGR when the generalization parameter is independent of spacetime quantities.
@@ Acknowledgment
Authors would like to express their gratitude to Monireh Emami for her valuable help in the numerical calculations.

### v2

@@ 4.2 The non-minimal coupling $q_{2}$
We now argue that the non-minimal coupling $q_{2}$ controls an effective scrambling time in the dual theory.
The non-minimal term $R^{2}F^{2}$ deforms Einstein-Maxwell theory and hence its holographic dual. Upon dimensional reduction, the undeformed theory is related to the Sachdev-Ye-Kitaev (SYK) model  [^5] We should note that the duality of SYK model and black holes (Jackiew-Teitelboim gravity) is well-established for the $q_{2}=0$, without the non-minimal coupling. With $q_{2}\neq 0$ the dual theory would be a deformed SYK model; nonetheless, at low-curvature regime it reduces to SYK model. In addition, our results about the curvature penalty factor is intact whether $q_{2}=0$ or not., which is known to saturate the bound on chaos and exhibit extremely fast scrambling. Introducing $q_{2}$ moves the system away from this maximally chaotic limit.
In the holographic description of quantum complexity, the growth rate of complexity is controlled not only by the bulk dynamics but also by the geometric structure induced by the complexity functional itself. In particular, within Nielsen’s geometric approach to circuit complexity, the scrambling time is understood as the timescale over which operators spread in the space of unitary transformations, a notion that depends on the choice of cost function and penalty factors.
In the present framework, the generalized complexity functional induces an effective geometry governing the optimal bulk trajectories associated with complexity growth. The presence of the generalized functional effectively rescales the metric entering the particle action (40) governing complexity  [^6] It is worth mentioning that this is a scaling and not the transformation of the coordinates.:
@@
where $g_{xx}$ and $g_{yy}$ are not scaled. As shown in (61), the generalized term rescales the metric entering the particle action that computes the complexity. Importantly, this rescaling does not alter the bulk equations of motion and therefore does not modify the intrinsic chaotic properties of the black brane background. Instead, it changes the notion of distance relevant for complexity, in direct analogy with the introduction of penalty factors in quantum circuit geometry.
As a consequence, the butterfly velocity  [^7] Butterfly velocity is a measure that how fast information, or the effect of a tiny perturbation, spreads through a quantum system. More precisely, for two local quantum operators $W$ and $V$ it is given by out-of-time-order correlators, $W(t,x),V(0)]^{2}>$, where at large $t$, its growth is proportional to $e^{\lambda_{L}(t-\frac{x}{v_{B}})}$ for the parameter $\lambda_{L}$, the Lyapunov exponent (how fast chaos spreads spatially). extracted from the effective metric differs from that associated with the physical spacetime metric. The effective butterfly velocity for the metric (61) is computed [44]:
@@
[…]
@@
This rescaling leads to an effective scrambling time (for details see Appendix B):
@@
It is important to emphasize that the scrambling time discussed in this section is not the microscopic scrambling time defined via out-of-time-order correlators or shock-wave analyses, which depends solely on the bulk equations of motion. Rather, we introduce an effective scrambling time that governs the rate of complexity growth. This timescale depends on the geometry entering the complexity functional and therefore can be sensitive to the generalized parameter defining the cost metric.
This interpretation is further supported by the near-horizon expansion of the CGR, where increasing $|q_{2}|$ generically leads to an enhancement of the effective scrambling time and a corresponding suppression of the CGR, in agreement with the numerical results presented below.
In the following it is shown that the CGR near-horizon expansion is inversely related to the scrambling time. The leading term in (64) would be:
@@
$\approx\frac{1}{t_{scal.}}=\frac{|a(R)|}{t_{*}}.$ (74)
Now, we get into the numerical results.
@@
The leading order in the CGR based on (73), (85), (86) and setting $\gamma=-0.002$ would be:
@@
Fig.s (5b) and (5d) illustrate that the CGR is reduced by the increase in the non-minimal coupling.
In this section we showed that the new scrambling time is related to the non-minimal coupling $q_{2}$ and consequently, by the change of the coupling, the CGR is changed as is expected.
@@ 4.3 The conserved charge $Q$
Finally, we interpret the conserved charge $Q$ and argue that it effectively restricts the set of simple unitary operators available in the dual quantum circuit. As a result, increasing $Q$ suppresses the complexity growth rate.
@@
In short, SKA states that for an arbitrary accuracy $\epsilon$, if $d(I,U)<\epsilon$, then there exist $V$ and $W$ such that $d(U,VWV^{\dagger}W^{\dagger})<\epsilon^{3/2}c_{1}$ and $d(I,V),d(I,W)<\sqrt{\epsilon}c_{2}$ where $I$ is identity operator, $c_{1}$ and $c_{2}$ are constant [45]. In addition, it can be shown [46] that if $V=e^{iF}$ and $W=e^{iG}$, then:
@@
These relations explicitly demonstrate that the size of $H||$ restricts the set of admissible generators for the simple unitary operators.
From the holographic perspective, the conserved charge $Q$ plays an analogous role. Increasing $Q$ effectively increases the scale associated with the generator $H$, thereby reducing the set of allowed simple unitaries in the boundary theory. This restriction manifests itself as a suppression of the rate at which complexity can grow.
This qualitative picture is supported by the bulk analysis. As discussed in the previous subsection, the effective scrambling time increases with $Q$. Since complexity growth is inversely related to the scrambling time, a larger charge leads to a smaller complexity growth rate. Explicitly, in the following subsections near the horizon one finds for all three choices of the generalization parameter that:
$\dot{\mathcal{C}}_{Any}\sim 1-\frac{Q^{2}}{2}+$..., (90)
indicating a decrease of the CGR as $Q$ increases. This behavior is in full agreement with the numerical results shown in Figs.(1), (3), and (5).
@@
The complexity growth rate based on (73), (75) and (76) is given by:
@@
Fig.s (3b) and (3d) displays that by increase in the conserved charge, the CGR is reduced.
@@
Fig.s (5b) and (5d) depict the decrease of the CGR by the increase in the conserved charge.
In summary, the conserved charge $Q$ restricts the available simple unitary operators in the dual circuit, thereby increasing the effective scrambling time and suppressing the complexity growth rate, in accordance with both the Solovay-Kitaev framework and the holographic analysis.
4.4 Superconducting qubits circuits
In this section, we provide two examples in superconducting qubits circuits. These examples demonstrate that modifying a penalty parameter can either suppress or enhance the effective processing rate, depending on how it constrains the circuit dynamics. This behavior closely parallels the dependence of the holographic complexity growth rate (CGR) on the generalization parameter $\gamma b$.
Example I: Crosstalk penalty and increased circuit depth. In the superconducting qubit circuits to prepare a qubit one needs controlling quickly the fast gates (or low runtime) to stay away from decoherence. However, fast controlling pose a problem because of the fact that it requires a strong, sharp microwave pulse, but strong pulse has a wide frequency spectrum. A strong pulse allows for very fast $CNOT$ gates [47]. Nonetheless, this makes parallelizing operations to reduce circuit depth. Moreover, a wide frequency pulse excite the qubit to higher energy states which leads to the leakage error or crosstalk between energy levels. To reduce the leakage it requires the qubit be more anharmonic that means the energy gap of the states be more significantly different [47]. The anharmonicity makes the states more isolated from the states with higher energies. A problem back of the other, the anharmonicity makes the qubit more sensitive to charge noise which results in reducing the coherence time. As a way out, one needs increase the circuit depth by applying crosstalk penalty.
@@ 4.4 Superconducting qubits circuits
Suppose that there are three qubits of QB1 connected to QB2 and QB2 connected to QB3 where crosstalk is high between QB1&\&QB2 and QB2&\&QB3. If one needs to apply a CNOT(1,3) operator, it is not possible due to the lack of direct coupling. Instead, one must first apply a SWAP(2,3) operation, followed by CNOT(1,3). Since, SWAP(2,3) gate decomposes into three CNOT gates, this procedure significantly increases the circuit depth, as illustrated in figure (7). Similar constraints arise in current IBM superconducting quantum processors [48].
In this setting, increasing the crosstalk penalty directly suppresses the processing rate. This behavior mirrors the holographic result in which increasing the penalty factor $\gamma b$ reduces the complexity growth rate.
@@
Example II: Penalty terms in quantum approximate optimization algorithms. A contrasting example is provided by quantum approximate optimization algorithms (QAOA), where penalty terms can be used to enhance computational efficiency rather than suppress it. In these algorithms, constraints are incorporated into the problem Hamiltonian as:
@@
where $\gamma_{a}$ controls the strength of the constraint (similar to $\gamma$ in the generalized complexity). Increasing $\gamma_{a}$ energetically penalizes states that violate the constraints, effectively guiding the system toward the desired subspace of solutions. As a result, the target state can be reached using fewer layers of quantum gates. In practice, this leads to a reduction in circuit depth, which is particularly advantageous in noisy superconducting devices where long circuits suffer from decoherence. Explicit examples of this behavior have been demonstrated in [49] where intensifying the obeying rules, the system in shorter steps gets to the desired solution.
Related optimization strategies have also been proposed in [50], known as mixer-phaser ansatz in superconducting processors in which they did not used a penalty-based algorithm. In a layman’s term, if the problem system goes away from your desired solution, the penalty factor penalizes the system to obey the constraints. However, in [50] the constraints have been inserted before the system circuit. Then, they get rid of the penalizing.
4.5 Relation to holographic complexity
These two examples illustrate that penalty factors in quantum circuits can have qualitatively different effects on the processing rate, depending on how they constrain the dynamics. In the holographic framework, this behavior is reflected in the dependence of the CGR on the choice of generalization parameter $b$. Certain choices lead to a suppression of the CGR, while others enhance it, in direct analogy with the superconducting-circuit examples discussed above.
Finally, we note that these circuit-based interpretations can be extended beyond qubit systems. SYK model, which describes strongly interacting fermions and serves as a holographic dual of near-$AdS_{2}$ gravity [51], can be mapped to qubit systems using the Bravyi-Kitaev [52] and Jordan-Wigner transformations [53]. This correspondence allows the circuit-based intuition developed here to be applied equally to strange metal systems.
@@ 5 Conclusion
In this work, we studied holographic complexity in four-dimensional Einstein-Maxwell theory with a non-minimal coupling of the form $R^{2}F_{\mu\nu}F^{\mu\nu}$. Since an exact black brane solution is not available, we constructed a perturbative AdS solution to first order in the non-minimal coupling parameter $q_{2}$. This model exhibits a linear temperature dependence of the resistivity and therefore provides a holographic realization of strange metal behavior, although no phase transition is present.
We analyzed the complexity of this system within the $complexity=anything$ framework, which allows for a generalized definition of holographic complexity. In this formulation, the complexity growth rate (CGR) depends on three independent parameters: the conserved charge $Q$, the non-minimal coupling $q_{2}$, and the choice of the generalized term $b$ entering the complexity functional. We explicitly considered three representative choices, namely the Weyl tensor squared $C^{2}$, $R^{2}F^{2}$, and the Maxwell invariant $F^{2}$.
We showed analytically that the generalized bulk functional deforms the effective geometry governing the extremal trajectories associated with complexity. This deformation modifies the effective cost metric in a manner structurally equivalent to the introduction of penalty factors in Nielsen’s geometric approach to quantum circuit complexity. Given the intrinsic ambiguity in the choice of gate sets and penalty schemes in holographic duality, this correspondence should be understood at the level of complexity geometry rather than as a microscopic identification of boundary gates. In this sense, the generalized parameter encodes a bulk realization of penalty structures that control the relative cost of different directions in complexity space.
We further examined the implications of this deformation for the CGR. By analyzing the effective butterfly velocity associated with the deformed complexity geometry, we identified an effective scrambling time that governs the CGR. We demonstrated analytically that this timescale depends on both the non-minimal coupling and the generalized parameter, while the intrinsic chaotic properties of the background geometry remain unchanged. Our numerical results support this analytical picture and show that increasing the strength of the generalized deformation, $|q_{2}|$ typically leads to a suppression of the CGR through an enhancement of the effective scrambling time.
Moreover, the conserved charge restricts the set of admissible simple unitary operators in the boundary theory, in accordance with the Solovay-Kitaev framework, and consequently suppresses the CGR.
To further support our interpretations, we discussed analogies with superconducting-qubit circuits, where penalty factors can either suppress or enhance the effective processing rate depending on how they constrain circuit dynamics. These examples demonstrate that the holographic behavior of the CGR is consistent with well-established mechanisms in realistic quantum computational platforms.
We found that the behavior of the CGR under variations of $q_{2}$ depends sensitively on the choice of the generalization parameter. For $b=C^{2}$ and $b=F^{2}$, increasing $q_{2}$ leads to a suppression of the CGR, whereas for $b=R^{2}F^{2}$ the opposite behavior is observed. Moreover, the generalization $b=F^{2}$ does not give rise to multiple branches in the CGR and closely resembles the behavior of the standard CV and CA proposals.
Several directions for future work are suggested by our results. It would be interesting to formulate a holographic complexity proposal that avoids explicit penalty factors and instead constrains the target state directly. Additionally, the appearance of multiple branches in the CGR for curvature-dependent generalizations, but not for spacetime-independent choices such as $F^{2}$, merits further investigation. Understanding this distinction may shed light on the geometric origin of complexity in holography.
@@ Acknowledgment
Authors would like to express their gratitude to Monireh Emami for her valuable help in the numerical calculations. Additionally, OpenAI’s ChatGPT was used to enhance the clarity of this manuscript.